# Universal Physics Transformers: A Framework For Efficiently Scaling Neural Operators

### v2

@@ 1 Introduction
In scientific pursuits, extensive efforts have produced highly intricate mathematical models of physical phenomena, many of which are naturally expressed as partial differential equations (PDEs) (Olver, 2014). Solving most PDEs is analytically intractable and necessitates falling back on compute-expensive numerical approximation schemes. In recent years, deep neural network based surrogates, most importantly neural operators (Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2021), have emerged as a computationally efficient alternative (Thuerey et al., 2021; Zhang et al., 2023), and impact e.g., weather forecasting (Lam et al., 2022; Bi et al., 2022; Andrychowicz et al., 2023), molecular modeling (Batzner et al., 2022; Batatia et al., 2022), or computational fluid dynamics (Vinuesa & Brunton, 2022; Guo et al., 2016; Li et al., 2020a; Kochkov et al., 2021; Gupta & Brandstetter, 2022; Carey et al., 2024). Additional to computational efficiency, neural surrogates offer potential to introduce generalization capabilities across phenomena, as well as generalization across characteristics such as boundary conditions or PDE coefficients (McCabe et al., 2023; Brandstetter et al., 2022b). Consequently, the nature of neural operators inherently complements handcrafted numerical solvers which are characterized by a substantial set of solver requirements, and mostly due to these requirements tend to differ among sub-problems (Bartels, 2016).
However, similar to their numerical counterparts, different neural network techniques are prevalent across applications. For example, when contrasting particle- and grid-based dynamics in computational fluid dynamics (CFD), i.e., Lagrangian and Eulerian discretization schemes. This is in contrast to other areas of deep learning where the flexibility of transformers (Vaswani et al., 2017) has enabled unified architectures across domains, allowing advancements in one domain to also benefit all others. This has lead to an efficient scaling of architectures, paving the way for large “foundation” models (Bommasani et al., 2021) that are pretrained on huge passive datasets  Devlin et al. (2018); He et al. (2022).
@@ 2 Background
In this work, we focus on experiments on (systems of) PDEs, that evolve a signal $\bm{u}(t,\bm{x})=\bm{u}^{t}(\bm{x})\in\mathbb{R}^{d}$ in a single temporal dimension $t\in[0,T]$ and $m$ spatial dimensions $\bm{x}\in U\subset\mathbb{R}^{m}$, for an open set $U$. With $1\leq l\in\mathbb{N}$, systems of PDEs of order $l$ can be written as
@@
We work mostly with the incompressible Navier-Stokes equations (Temam, 2001), which e.g., in two spatial dimensions conserve the velocity flow field $\bm{u}(t,x,y):[0,T]\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ via:
@@
where $\bm{u}\cdot\nabla\bm{u}$ is the convection, i.e., the rate of change of $\bm{u}$ along $\bm{u}$, $\mu$ is the viscosity parameter, $\mu\nabla^{2}\bm{u}$ the viscosity, i.e., the diffusion or net movement of $\bm{u}$, $\nabla p$ the internal pressure gradient, and $\bm{f}$ an external force. The constraint $\nabla\cdot\bm{u}=0$ yields mass conservation of the Navier-Stokes equations. A detailed depiction of the involved differential operators is given in Appendix A.1.
@@
Operator learning (Lu et al., 2019, 2021; Li et al., 2020b, a; Kovachki et al., 2021) learns a mapping between function spaces, and this concept is often used to approximate solutions of PDEs. Similar to Kovachki et al. (2021), we assume $\mathcal{U},\mathcal{V}$ to be Banach spaces of functions on compact domains $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ or $\mathcal{Y}\subset\mathbb{R}^{d_{y}}$, mapping into $\mathbb{R}^{d_{u}}$ or $\mathbb{R}^{d_{v}}$, respectively. The goal of operator learning is to learn a ground truth operator ${\mathcal{G}}:\mathcal{U}\rightarrow\mathcal{V}$ via an approximation $\hat{{\mathcal{G}}}:\mathcal{U}\rightarrow\mathcal{V}$. This is usually done in the vein of supervised learning by i.i.d. sampling input-output pairs, with the notable difference, that in operator learning the spaces sampled from are not finite dimensional. More precisely, with a given data set consisting of $N$ function pairs $(\bm{u}_{i},\bm{v}_{i})=(\bm{u}_{i},{\mathcal{G}}(\bm{u}_{i}))\subset\mathcal{U}\times\mathcal{V}$, $i=1,...N$, we aim to learn $\hat{{\mathcal{G}}}:\mathcal{U}\rightarrow\mathcal{V}$, so that ${\mathcal{G}}$ can be approximated in a suitably chosen norm.
@@
A popular approach, that we are also going to follow, is to approximate ${\mathcal{G}}$ via three maps (Seidman et al., 2022): ${\mathcal{G}}\approx\hat{{\mathcal{G}}}\coloneqq\mathcal{D}\circ\mathcal{A}\circ\ \mathcal{E}.$ The encoder $\mathcal{E}:\mathcal{U}\rightarrow\mathbb{R}^{h_{1}}$ takes an input function and maps it to a finite dimensional latent feature representation. For example, $\mathcal{E}$ could embed a continuous function to a chosen hidden dimension $\mathbb{R}^{h_{1}}$ for a collection of grid points. Next, $\mathcal{A}:\mathbb{R}^{h_{1}}\rightarrow\mathbb{R}^{h_{2}}$ approximates the action of the operator $\mathcal{G}$, and $\mathcal{D}$ decodes the hidden representation, and thus creates the output functions via $\mathcal{D}:\mathbb{R}^{h_{2}}\rightarrow\mathcal{V}$, which in many cases is point-wise evaluated at the output grid or output mesh.
@@
Often, numerical simulation methods can be classified into two distinct families: particle and grid-based methods. This specification is notably prevalent, for instance, in the field of computational fluid dynamics (CFD), where Lagrangian and Eulerian discretization schemes offer different characteristics dependent on the PDEs. In simpler terms, Eulerian schemes essentially monitor velocities at specific fixed grid points. These points, represented by a spatially limited number of nodes, control volumes, or cells, serve to discretize the continuous space. This process leads to grid-based or mesh-based representations. In contrast to such grid- and mesh-based representations, in Lagrangian schemes, the discretization is carried out using finitely many material points, often referred to as particles, which move with the local deformation of the continuum. Roughly speaking, there are three families of Lagrangian schemes: discrete element methods (Cundall & Strack, 1979), material point methods (Sulsky et al., 1994; Brackbill & Ruppel, 1986), and smoothed particle hydrodynamics (SPH) (Gingold & Monaghan, 1977; Lucy, 1977; Monaghan, 1992, 2005). In this work, we focus on SPH methods, which approximate the field properties using radial kernel interpolations over adjacent particles at the location of each particle. The strength of SPH lies in its ability to operate without being constrained by connectivity issues, such as meshes. This characteristic proves especially beneficial when simulating systems that undergo significant deformations.
@@
For larger meshes or larger number of particles, memory consumption and inference speed become more and more important. Fourier Neural Operator (FNO) based methods work on regular grids, or learn a mapping to a regular latent grid, e.g., geometry-informed neural operators (GINO) (Li et al., 2023). In three dimensions, the stored Fourier modes have the shape $h\times n_{x}\times n_{y}\times n_{z}$, where $h$ is the hidden size and $n_{x}$, $n_{y}$, $n_{z}$ are the respective Fourier modes. Similarly, the latent space of CNN-based methods, e.g., Raonić et al. (2023); Gupta & Brandstetter (2022), is of shape $h\times w_{x}\times w_{y}\times w_{z}$, where $w_{x}$, $w_{y}$, $w_{z}$ are the respective grid points. In three dimension, the memory requirement in each layer increases cubically with increasing number of modes or grid points. In contrast, transformer based neural operators, e.g., Hao et al. (2023); Cao (2021); Li et al. (2022), operate on a token-based latent space of dimension $n_{\text{tokens}}\times h$, where usually $n_{\text{tokens}}\propto n_{\text{points}}$, and GNN based neural operators, e.g., Li et al. (2020b), operate on a node based latent space of dimension $n_{\text{nodes}}\times h$, where usually $n_{\text{nodes}}=n_{\text{points}}$. For large number of inputs, this becomes infeasible as every layer has to process a large number of tokens. Contrary, UPTs compress the inputs into a low-dimensional latent space, which drastically decreases computational requirements. Different architectures are compared in Tab. 1. As it can be hard to imagine the practical relevance of theoretical complexity measures, we qualitatively study the scaling limits of representative models in Fig. 2 where we find the efficient latent space compression of UPTs allows training on up to 4.2M input points which is 64x the amount that a GNN could handle.
Model Range Complexity Irregular Discretization Learns Latent
Grid Convergent Field Rollout
GNN local $O(MD)$
CNN local $O(G)$
Transformer global $O(M^{2})$
Linear Transformer global $O(M)$
GNO (Li et al., 2020b) radius $O(MD)$
FNO (Li et al., 2020a) global $O(G\log G)$
GINO (Li et al., 2023) global $O(GD+G\log G)$
UPT (with GNO) global $O(SD+S^{2})$
UPT (without GNO) global $O(SM+S^{2})$
Table 1: Model comparison. Complexity includes number of mesh points $M$, and maximum degree of the graph $D$. Grid-based methods project the mesh to $G$ grid points. UPTs instead use a small amount of supernodes $S$ as discretization, where $G$ is typically much larger than $S$. The UPT training procedure separates responsibilities between components, allowing us to forward propage dynamics purely within the latent space. For large meshes, UPT uses a GNO to compress the mesh points into supernodes (“with GNO”), which can be omitted for small meshes (“without GNO”).
Figure 2: Qualitative exploration of scaling limits. Starting from 32K input points (scale 1), we train a 68M parameter model for a few steps with batchsize $1$ and measure the required GPU memory. “LT” denotes a transformer with linear attention. Models without a compressed latent space (GNN, LT) quickly reach their limits while models with a compressed latent space (GINO, UPT) scale much better with the number of inputs. However, as GINO compresses the latent space onto a regular grid, the scaling benefits are largely voided on 3D problems. The efficient latent space compression of UPTs can fit up to 4.2M points (scale 128). Implementation details are outlined in App. B.7.
@@ 3 Universal Physics Transformers
Our goal is to learn a mapping between the solutions $\bm{u}^{t}$ and $\bm{u}^{t^{\prime}}$ of Eq. (1) at timesteps $t$ and $t^{\prime}$, respectively. Our dataset should consist of $N$ function pairs $(\bm{u}^{t}_{i},\bm{u}^{t^{\prime}}_{i})$, $i=1,..,N$, where each $\bm{u}^{t}_{i}$ is sampled at $k$ spatial locations $\{\bm{x}^{1}_{i},\ldots,\bm{x}^{k}_{i}\}\in U$. Similarly, we query each output signal $\hat{\bm{u}}^{t^{\prime}}_{i}$ at $k^{\prime}$ spatial locations $\{\bm{y}^{1}_{i},\ldots,\bm{y}^{k^{\prime}}_{i}\}\in U$. Then each input signal can be represented by $\bm{u}_{i,k}^{t}=(\bm{u}_{i}^{t}(\bm{x}^{1}_{i}),\ldots,\bm{u}_{i}^{t}(\bm{x}^{k}_{i}))^{T}\in\mathbb{R}^{k\times d}$ as a tensor of shape $k\times d$, similar for the output. For particle- or mesh-based inputs, it is often simpler to represent the input as graph $\mathscr{G}=(V,E)$ with $k$ nodes $\{\bm{x}^{1}_{i},\ldots\bm{x}^{k}_{i}\}\in V$, edges $E$ (that reflect the neighborhood structure) and node features $\{\bm{u}^{t}_{i}(\bm{x}^{1}_{i}),\ldots,\bm{u}^{t}_{i}(\bm{x}^{k}_{i})\}$.
Figure 3: Schemantic sketch of the UPT architecture. The encoder compresses information from various grids or differing particles. Subsequently, this information is propagated forward in time through the approximator and decoded at arbitrary query positions.
@@
The encoder $\mathcal{E}$ first embeds $k$ points into hidden dimension $h$, adding position encoding (Vaswani et al., 2017) to the different nodes, i.e., $\bm{u}_{i,k}^{t}\in\mathbb{R}^{k\times d}\rightarrow\mathbb{R}^{k\times h}$. In the first hierarchy, information is exchanged between local points and a selected set of $n_{\text{s}}$ supernode points. For Eulerian discretization schemes those supernodes can either be uniformly sampled on a regular grid as in Li et al. (2023), or selected based on the given mesh. The latter has the advantage that mesh characteristics are automatically taken into account, e.g., dense or sparse mesh regions are represented by different numbers of nodes. Furthermore, adaptation to new meshes is straightforward. We implement the first hierarchy by randomly selecting $n_{\text{s}}$ supernodes on the mesh, choosing $n_{\text{s}}$ such that the mesh characteristic is preserved. E.g., for experiments presented in Sec. 4.2, $n_{\text{s}}=2048$. Similarly, in the Lagrangian discretization scheme, choosing supernodes based on particle positions provides the same advantages as selecting them based on the mesh.
Information is aggregated at the selected $n_{\text{s}}$ supernodes via a message passing layer (Gilmer et al., 2017) using a radius graph between points. Importantly, messages only flow towards the $n_{\text{s}}$ supernodes, and thus the compute complexity of the first hierarchy scales linearly with $n_{\text{s}}$. The second hierarchy consists of transformer blocks (Vaswani et al., 2017) followed by a perceiver block (Jaegle et al., 2021b, a) with $n_{\text{latent}}$ learned queries of dimension $h$. To summarize, the encoder $\mathcal{E}$ maps $\bm{u}_{i}^{t}\in\mathcal{U}$ to a latent space via
@@
where tyically $n_{\text{latent}}\ll n_{\text{s}}<k$. If the number of points is manageable, the first hierarchy can be omitted.
@@
The approximator propagates the compressed representation forward in time. As $n_{\text{latent}}$ is small, forward propagation in time is fast. We employ a stack of transformer blocks as approximator.
@@
Notably, the approximator can be applied multiple times, propagating the input signal forward in time by $\Delta t$ each time. If $\Delta t$ is small enough, the input signal can be approximated at arbitrary future times $t^{\prime}$.
@@
The decoder is implemented via a perceiver-like cross attention layer using a positional embedding of the output positions as query and the latent representation $\bm{z}_{i}^{t^{\prime}}$ as keys and values. Since there is no interaction between queries, the latent representation can be queried at arbitrarily many positions without large computational overhead. This decoding mechanism establishes a connection of conditioned neural fields to operator learning (Perdikaris, 2023).
@@
To condition the model to the current timestep $t$ and to boundary conditions such as the inflow velocity, we add feature modulation to all transformer and perceiver blocks. We use DiT modulation (Peebles & Xie, 2023), which consists of a dimension-wise scale, shift and gate operation that are applied to the attention and MLP module of the transformer. Scale, shift and gate are dependent on an embedding of the timestep and boundary conditions (e.g. velocity).
@@
Figure 4: UPT training procedure to enable latent rollouts. To separate the responsibilities of the individual components, we introduce inverse encoding and decoding losses in addition to the next step prediction loss.
@@
The closest work to ours are transformer neural operators of Cao (2021); Li et al. (2022); Hao et al. (2023) which encode different query points into a tokenized latent space representation of dimension $n_{\text{nodes}}\times h$, where $n_{\text{nodes}}$ varies based on the number of input points, i.e., $n_{\text{nodes}}\propto n_{\text{points}}$. Wu et al. (2024) adds a learnable mapping into a fixed latent space of dimension $n_{\text{nodes}}\times h$ to each transformer layer, and projects back to dimension $n_{\text{points}}\times h$ after self-attention. In contrast, UPTs use fixed $n_{\text{latent}}$ for the unified latent space representation $n_{\text{latent}}\times h$.
For the modeling of temporal PDEs, a common scheme is to map the input solution at time $t$ to the solution at next time step $t^{\prime}$ (Li et al., 2020a; Brandstetter et al., 2022a; Takamoto et al., 2022). Especially for systems that are modeled by graph-based representations, predicted accelerations at nodes are numerically integrated to model the time evolution of the system (Sanchez-Gonzalez et al., 2020; Pfaff et al., 2020). Recently, equivariant graph neural operators (Xu et al., 2024) were introduced which model time evolution via temporal convolutions in Fourier space. More related to our work are methods that propagate dynamics in the latent space (Lee & Carlberg, 2021; Wiewel et al., 2019). Once the system is encoded, time evolution is modeled via LSTMs (Wiewel et al., 2019), or even linear propagators (Lusch et al., 2018; Morton et al., 2018). In Li et al. (2022), attention-based layers are used for encoding the spatial information of the input and query points, while time updates in the latent space are performed using recurrent MLPs. Similarly, Bryutkin et al. (2024) use recurrent MLPs for temporal updates within the latent space, while utilizing a graph transformer for encoding the input observations.
Building universal models aligns with the contemporary trend of foundation models for science. Recent works comprise pretraining over multiple heterogeneous physical systems, mostly in the form of PDEs (McCabe et al., 2023), foundation models for weather and climate (Nguyen et al., 2023), or material modeling (Merchant et al., 2023; Zeni et al., 2023; Batatia et al., 2022).
Methods similar to our latent space modeling have been proposed in the context of diffusion models (Rombach et al., 2022) where a pre-trained compression model is used to compress the input into a latent space from which a diffusion model can be trained at much lower costs. Similarly, our approach also compresses the high-dimensional input into a low-dimensional latent space, but without relying on a two stage approach. Instead, we learn the compression end-to-end via inverse encoding and decoding techniques.

### v3

@@ 1 Introduction
In scientific pursuits, extensive efforts have produced highly intricate mathematical models of physical phenomena, many of which are naturally expressed as partial differential equations (PDEs) [76]. Solving most PDEs is analytically intractable and necessitates falling back on compute-expensive numerical approximation schemes. In recent years, deep neural network based surrogates, most importantly neural operators [51, 61, 47], have emerged as a computationally efficient alternative [99, 119], and impact e.g., weather forecasting [48, 8, 1], molecular modeling [7, 4], or computational fluid dynamics [107, 30, 51, 43, 31, 19]. Additional to computational efficiency, neural surrogates offer potential to introduce generalization capabilities across phenomena, as well as generalization across characteristics such as boundary conditions or PDE coefficients [66, 14]. Consequently, the nature of neural operators inherently complements handcrafted numerical solvers which are characterized by a substantial set of solver requirements, and mostly due to these requirements tend to differ among sub-problems [3].
However, similar to their numerical counterparts, different neural network techniques are prevalent across applications. For example, when contrasting particle- and grid-based dynamics in computational fluid dynamics (CFD), i.e., Lagrangian and Eulerian discretization schemes. This is in contrast to other areas of deep learning where the flexibility of transformers [106] has enabled unified architectures across domains, allowing advancements in one domain to also benefit all others. This has lead to an efficient scaling of architectures, paving the way for large “foundation” models [10] that are pretrained on huge passive datasets  [25, 36].
@@ 2 Background
We focus on experiments on (systems of) PDEs, that evolve a signal $\bm{u}(t,\bm{x})=\bm{u}^{t}(\bm{x})\in\mathbb{R}^{d}$ in a single temporal dimension $t\in[0,T]$ and $m$ spatial dimensions $\bm{x}\in U\subset\mathbb{R}^{m}$, for an open set $U$. With $1\leq l\in\mathbb{N}$, systems of PDEs of order $l$ can be written as
@@
We work mostly with the incompressible Navier-Stokes equations [98], which e.g., in two spatial dimensions conserve the velocity flow field $\bm{u}(t,x,y):[0,T]\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ via:
@@
where $\bm{u}\cdot\nabla\bm{u}$ is the convection, i.e., the rate of change of $\bm{u}$ along $\bm{u}$, $\mu$ is the viscosity parameter, $\mu\nabla^{2}\bm{u}$ the viscosity, i.e., the diffusion or net movement of $\bm{u}$, $\nabla p$ the internal pressure gradient, and $\bm{f}$ an external force. The constraint $\nabla\cdot\bm{u}=0$ yields mass conservation of the Navier-Stokes equations. A detailed depiction of the involved differential operators is given in Appendix B.1.
@@
Operator learning [60, 61, 52, 51, 47] learns a mapping between function spaces – a concept which is often used to approximate solutions of PDEs. Similar to Kovachki et al. [47], we assume $\mathcal{U},\mathcal{V}$ to be Banach spaces of functions on compact domains $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ or $\mathcal{Y}\subset\mathbb{R}^{d_{y}}$, mapping into $\mathbb{R}^{d_{u}}$ or $\mathbb{R}^{d_{v}}$, respectively. The goal of operator learning is to learn a ground truth operator ${\mathcal{G}}:\mathcal{U}\rightarrow\mathcal{V}$ via an approximation $\hat{{\mathcal{G}}}:\mathcal{U}\rightarrow\mathcal{V}$. This is usually done in the vein of supervised learning by i.i.d. sampling input-output pairs, with the notable difference, that in operator learning the spaces sampled from are not finite dimensional. More precisely, with a given data set consisting of $N$ function pairs $(\bm{u}_{i},\bm{v}_{i})=(\bm{u}_{i},{\mathcal{G}}(\bm{u}_{i}))\subset\mathcal{U}\times\mathcal{V}$, $i=1,...N$, we aim to learn $\hat{{\mathcal{G}}}:\mathcal{U}\rightarrow\mathcal{V}$, so that ${\mathcal{G}}$ can be approximated in a suitably chosen norm.
@@
We follow the popular approach to approximate ${\mathcal{G}}$ via three maps [89]: ${\mathcal{G}}\approx\hat{{\mathcal{G}}}\coloneqq\mathcal{D}\circ\mathcal{A}\circ\ \mathcal{E}.$ The encoder $\mathcal{E}:\mathcal{U}\rightarrow\mathbb{R}^{h_{1}}$ takes an input function and maps it to a finite dimensional latent feature representation. For example, $\mathcal{E}$ could embed a continuous function to a chosen hidden dimension $\mathbb{R}^{h_{1}}$ for a collection of grid points. Next, $\mathcal{A}:\mathbb{R}^{h_{1}}\rightarrow\mathbb{R}^{h_{2}}$ approximates the action of the operator $\mathcal{G}$, and $\mathcal{D}$ decodes the hidden representation, and thus creates the output functions via $\mathcal{D}:\mathbb{R}^{h_{2}}\rightarrow\mathcal{V}$, which in many cases is point-wise evaluated at the output grid or output mesh.
@@
Often, numerical simulation methods can be classified into two distinct families: particle and grid-based methods. This specification is notably prevalent, for instance, in the field of computational fluid dynamics (CFD), where Lagrangian and Eulerian discretization schemes offer different characteristics dependent on the PDEs. In simpler terms, Eulerian schemes essentially monitor velocities at specific fixed grid points. These points, represented by a spatially limited number of nodes, control volumes, or cells, serve to discretize the continuous space. This process leads to grid-based or mesh-based representations. In contrast to such grid- and mesh-based representations, in Lagrangian schemes, the discretization is carried out using finitely many material points, often referred to as particles, which move with the local deformation of the continuum. Roughly speaking, there are three families of Lagrangian schemes: discrete element methods [24], material point methods [94, 12], and smoothed particle hydrodynamics (SPH) [29, 62, 69, 70]. In this work, we focus on SPH methods, which approximate the field properties using radial kernel interpolations over adjacent particles at the location of each particle. The strength of SPH lies in its ability to operate without being constrained by connectivity issues, such as meshes. This characteristic proves especially beneficial when simulating systems that undergo significant deformations.
Model Range Irregular Discr. Learns Latent
Grid Conv. Field Rollout
GNN local
CNN local
Transformer global
GNO [52] radius
FNO [51] global
GINO [54] global
UPT global
Figure 2: Qualitative exploration of scaling limits. Starting from 32K input points (scale 1), we train a 68M parameter model for a few steps with batchsize $1$ and measure the required GPU memory. Models without a compressed latent space (GNN, Transformer) quickly reach their limits while models with a compressed latent space (GINO, UPT) scale much better with the number of inputs. However, as GINO compresses the latent space onto a regular grid, the scaling benefits are largely voided on 3D problems. The efficient latent space compression of UPTs can fit up to 4.2M points (scale 128). We use a linear attention transformer [90] for this study. “Disc. Conv.” denotes “Discretization Convergent”. Appendix D.7 outlines implementation details and complexities.
@@
For larger meshes or larger number of particles, memory consumption and inference speed become more and more important. Fourier Neural Operator (FNO) based methods work on regular grids, or learn a mapping to a regular latent grid, e.g., geometry-informed neural operators (GINO) [54]. In three dimensions, the stored Fourier modes have the shape $h\times n_{x}\times n_{y}\times n_{z}$, where $h$ is the hidden size and $n_{x}$, $n_{y}$, $n_{z}$ are the respective Fourier modes. Similarly, the latent space of CNN-based methods, e.g., Raonić et al. [83], Gupta & Brandstetter [31], is of shape $h\times w_{x}\times w_{y}\times w_{z}$, where $w_{x}$, $w_{y}$, $w_{z}$ are the respective grid points. In three dimension, the memory requirement in each layer increases cubically with increasing number of modes or grid points. In contrast, transformer based neural operators, e.g., Hao et al. [33], Cao [18], Li et al. [53], operate on a token-based latent space of dimension $n_{\text{tokens}}\times h$, where usually $n_{\text{tokens}}\propto n_{\text{points}}$, and GNN based neural operators, e.g., Li et al. [52], operate on a node based latent space of dimension $n_{\text{nodes}}\times h$, where usually $n_{\text{nodes}}=n_{\text{points}}$. For large number of inputs, this becomes infeasible as every layer has to process a large number of tokens. Contrary, UPTs compress the inputs into a low-dimensional latent space, which drastically decreases computational requirements. Different architectures and their scaling limits are compared in Fig. 2.
@@ 3 Universal Physics Transformers
Our goal is to learn a mapping between the solutions $\bm{u}^{t}$ and $\bm{u}^{t^{\prime}}$ of Eq. (1) at timesteps $t$ and $t^{\prime}$, respectively. Our dataset should consist of $N$ function pairs $(\bm{u}^{t}_{i},\bm{u}^{t^{\prime}}_{i})$, $i=1,..,N$, where each $\bm{u}^{t}_{i}$ is sampled at $k$ spatial locations $\{\bm{x}^{1}_{i},\ldots,\bm{x}^{k}_{i}\}\in U$. Similarly, we query each output signal $\hat{\bm{u}}^{t^{\prime}}_{i}$ at $k^{\prime}$ spatial locations $\{\bm{y}^{1}_{i},\ldots,\bm{y}^{k^{\prime}}_{i}\}\in U$. Then each input signal can be represented by $\bm{u}_{i,k}^{t}=(\bm{u}_{i}^{t}(\bm{x}^{1}_{i}),\ldots,\bm{u}_{i}^{t}(\bm{x}^{k}_{i}))^{T}\in\mathbb{R}^{k\times d}$ as a tensor of shape $k\times d$, similar for the output. For particle- or mesh-based inputs, it is often simpler to represent the input as graph $\mathscr{G}=(V,E)$ with $k$ nodes $\{\bm{x}^{1}_{i},\ldots\bm{x}^{k}_{i}\}\in V$, edges $E$ and node features $\{\bm{u}^{t}_{i}(\bm{x}^{1}_{i}),\ldots,\bm{u}^{t}_{i}(\bm{x}^{k}_{i})\}$.
Figure 3: Left: UPT compresses information from various grids or differing particles with an encoder, propagates this information forward in time through the approximator and decodes at arbitrary query positions. Right: Training procedure to enable latent rollouts via inverse encoding/decoding losses.
@@
The encoder $\mathcal{E}$ first embeds $k$ points into hidden dimension $h$, adding position encoding [106] to the different nodes, i.e., $\bm{u}_{i,k}^{t}\in\mathbb{R}^{k\times d}\rightarrow\mathbb{R}^{k\times h}$. In the first hierarchy, information is exchanged between local points and a selected set of $n_{\text{s}}$ supernode points. For Eulerian discretization schemes those supernodes can either be uniformly sampled on a regular grid as in [54], or selected based on the given mesh. The latter has the advantage that mesh characteristics are automatically taken into account, e.g., dense or sparse mesh regions are represented by different numbers of nodes. Furthermore, adaptation to new meshes is straightforward. We implement the first hierarchy by randomly selecting $n_{\text{s}}$ supernodes on the mesh, choosing $n_{\text{s}}$ such that the mesh characteristic is preserved. Similarly, in the Lagrangian discretization scheme, choosing supernodes based on particle positions provides the same advantages as selecting them based on the mesh.
Information is aggregated at the selected $n_{\text{s}}$ supernodes via a message passing layer [28] using a radius graph between points. Importantly, messages only flow towards the $n_{\text{s}}$ supernodes, and thus the compute complexity of the first hierarchy scales linearly with $n_{\text{s}}$. The second hierarchy consists of transformer blocks [106] followed by a perceiver block [40, 39] with $n_{\text{latent}}$ learned queries of dimension $h$. To summarize, the encoder $\mathcal{E}$ maps $\bm{u}_{i}^{t}\in\mathcal{U}$ to a latent space via
@@
where tyically $n_{\text{latent}}\ll n_{\text{s}}<k$. If the number of points is small, the first hierarchy can be omitted.
Note that randomly sampling mesh cells or particles implicitly encodes the underlying mesh or particle density and allocates more supernodes to highly resolved areas in the mesh or densely populated particle regions. Therefore, this can be seen as an implicit “importance sampling” of the underlying simulation. Additional implementation details are provided in Appendix F.1.
This encoder design projects into a fixed size latent space as it is an efficient way to compress the input into a fixed size representation to enable scaling to large-scale systems while remaining compute efficient. However, if an application requires a variable sized latent space, one could also remove the perceiver pooling layer. With this change the number of supernodes is equal to the number of latent tokens and complex problems could be tackled by a larger supernode count.
@@
The approximator propagates the compressed representation forward in time. As $n_{\text{latent}}$ is small, forward propagation in time is fast. We employ a transformer as approximator.
@@
Notably, the approximator can be applied multiple times, propagating the signal forward in time by $\Delta t$ each time. If $\Delta t$ is small enough, the input signal can be approximated at arbitrary future times $t^{\prime}$.
@@
The decoder is implemented via a perceiver-like cross attention layer using a positional embedding of the output positions as query and the latent representation $\bm{z}_{i}^{t^{\prime}}$ as keys and values. Since there is no interaction between queries, the latent representation can be queried at arbitrarily many positions without large computational overhead. This decoding mechanism establishes a connection of conditioned neural fields to operator learning [79].
@@
To condition the model to the current timestep $t$ and to boundary conditions such as the inflow velocity, we add feature modulation to all transformer and perceiver blocks. We use DiT modulation [78], which consists of a dimension-wise scale, shift and gate operation that are applied to the attention and MLP module of the transformer. Scale, shift and gate are dependent on an embedding of the timestep and boundary conditions (e.g. velocity).
@@
The closest work to ours are transformer neural operators of Cao [18], Li et al. [53], Hao et al. [33] which encode different query points into a tokenized latent space representation of dimension $n_{\text{nodes}}\times h$, where $n_{\text{nodes}}$ varies based on the number of input points, i.e., $n_{\text{nodes}}\propto n_{\text{points}}$. Wu et al. [114] adds a learnable mapping into a fixed latent space of dimension $n_{\text{nodes}}\times h$ to each transformer layer, and projects back to dimension $n_{\text{points}}\times h$ after self-attention. In contrast, UPTs use fixed $n_{\text{latent}}$ for the unified latent space representation $n_{\text{latent}}\times h$.
For the modeling of temporal PDEs, a common scheme is to map the input solution at time $t$ to the solution at next time step $t^{\prime}$ [51, 13, 95]. Especially for systems that are modeled by graph-based representations, predicted accelerations at nodes are numerically integrated to model the time evolution of the system [87, 81]. Recently, equivariant graph neural operators [116] were introduced which model time evolution via temporal convolutions in Fourier space. More related to our work are methods that propagate dynamics in the latent space [50, 111]. Once the system is encoded, time evolution is modeled via LSTMs [111], or even linear propagators [63, 73]. In Li et al. [53], attention-based layers are used for encoding the spatial information of the input and query points, while time updates in the latent space are performed using recurrent MLPs. Similarly, Bryutkin et al. [16] use recurrent MLPs for temporal updates within the latent space, while utilizing a graph transformer for encoding the input observations.
Building universal models aligns with the contemporary trend of foundation models for science. Recent works comprise pretraining over multiple heterogeneous physical systems, mostly in the form of PDEs [66], foundation models for weather and climate [75], or material modeling [67, 118, 4].
Methods similar to our latent space modeling have been proposed in the context of diffusion models [85] where a pre-trained compression model is used to compress the input into a latent space from which a diffusion model can be trained at much lower costs. Similarly, our approach also compresses the high-dimensional input into a low-dimensional latent space, but without a two stage approach. Instead, we learn the compression end-to-end via inverse encoding and decoding techniques.